Solve the inequality, and write the solution set in interval notation if possible. Write numbers as simplified fractions or integers.

|8x+1| +8<3

The solution set is....
Answer #1

There is no solution for the inequality.

Because |8x+1| is always non-negative .

That is |8x+1| >=0

So , |8x+1| + 8 >= 0+8 = 8.

The value of |8x+1| +8 is not less than 8

The given inequality is |8x+1| + 8 < 3 which is impossible because the Left hand side can not be less than 3 .

So there is no solution . The solution set is null set